# Combining Deterministic Enhanced Conditions with Dual-Streaming Encoding for Diffusion-Based Speech Enhancement

### v1

@@ Abstract
Diffusion-based speech enhancement (SE) models need to incorporate correct prior knowledge as reliable conditions to generate accurate predictions. However, providing reliable conditions using noisy features is challenging. One solution is to use features enhanced by deterministic methods as conditions. However, the information distortion and loss caused by deterministic methods might affect the diffusion process. In this paper, we first investigate the effects of using different deterministic SE models as conditions for diffusion. We validate two conditions depending on whether the noisy feature was used as part of the condition: one using only the deterministic feature (deterministic-only), and the other using both deterministic and noisy features (deterministic-noisy). Preliminary investigation found that using deterministic enhanced conditions improves hearing experiences on real data, while the choice between using deterministic-only or deterministic-noisy conditions depends on the deterministic models. Based on these findings, we propose a dual-streaming encoding Repair-Diffusion Model for SE (DERDM-SE) to more effectively utilize both conditions. Moreover, we found that fine-grained deterministic models have greater potential in objective evaluation metrics, while UNet-based deterministic models provide more stable diffusion performance. Therefore, in the DERDM-SE, we propose a deterministic model that combines coarse- and fine-grained processing. Experimental results on CHiME4 show that the proposed models effectively leverage deterministic models to achieve better SE evaluation scores, along with more stable performance compared to other diffusion-based SE models.
Index Terms:
Diffusion model, deterministic method, probabilistic method, ensemble, Brownian motion trajectory
@@ I Introduction
Speech enhancement (SE) is commonly used to recover speech components from noisy speech [1, 2, 3, 4, 5, 6, 7] to improve the speech quality and intelligibility. Supervised SE methods [8] generally outperform traditional SE methods [1, 9] with sufficient training data available. These systems can be broadly categorized as either deterministic [10, 11, 12, 13, 14] or probabilistic [15, 16, 17, 18, 19, 20, 21] approaches. Deterministic SE models learn an optimal deterministic mapping from noisy speech to clean speech [17], while probabilistic SE models capture the target distribution in either implicit or explicit way [15, 16, 17, 18]. The presence of diverse and uncontrollable sources often makes the noise unpredictable, resulting in random signals. However, deterministic SE models only learn a deterministic mapping function, leading to performance degradation in unseen noise conditions.
@@
Diffusion model [22, 23, 17, 16, 20, 24] is a probabilistic method learning a partially stochastic function. It contains a forward diffusion process and a reverse diffusion process by following the stochastic differential equation (SDE). The forward diffusion process gradually transforms the clean speech feature into noise, during which the neural network learns to reverse the incremental process of noise addition. The reverse process recovers the speech feature from noise with numerous reverse iterations. Different from the deterministic models [10, 11, 12, 13], SDE has the characteristics of the partially stochastic function as: random noise is injected into each reverse SDE step. Currently, diffusion model for SE [16, 17, 25] takes clean speech feature as the initial state and transforms it to noisy speech feature with Gaussian white noise. Noisy speech feature is used as condition [16, 17, 25] and drift term [17, 25] to enable the model to capture noise information implicitly. However, noisy features pose a challenge in providing accurate prior knowledge as reliable conditions for diffusion.
@@
Combining deterministic and diffusion models [26, 25, 19] makes a promising solution for alleviating this challenge. Both noisy and deterministically enhanced (deterministic-noisy) features or only deterministically enhanced (deterministic-only) features are used as conditions for diffusion [25]. When using a diffusion model to process deterministic enhanced features, the posterior samples are more likely to lie in high-density regions of the posterior probability for clean speech features given noisy speech features[25]. The effectiveness of the combination has been demonstrated [25, 26], but no previous work has investigated how different deterministic models affect the SE performance based on this pipeline.
@@
In this paper, we investigate the effects of using different deterministic SE models [27, 11, 28, 29] as conditions for diffusion-based SE. We experiment with diffusion models that use either deterministic-only enhanced features (Deter-Diffusion) or deterministic-noisy features (StoRM [25] and GP-Unified [19]) as conditions. Deter-Diffusion and StoRM train the diffusion model using only the score error, without repairing deterministic information. Compared to using noisy features as conditions, using deterministic enhanced features provides considerable improvements in hearing experiences [25]. However, the lack of deterministic information repair leads to instability in the performance of diffusion model. This is reflected in the fact that deterministic SE models with better objective evaluations do not lead to better diffusion performance. In addition to the score error, GP-Unified [19] introduces deterministic information into the diffusion model using an additional decoder. Although GP-Unified [19] outperforms StoRM [25], performing both deterministic estimation and diffusion within a shared bottleneck processing module may lead to greater performance degradation under noise mismatch conditions [24], as both the deterministic and score-estimation information exhibit mismatched data distributions between simulated and real data.
@@
To achieve more stable diffusion performance, we propose a Repair-Diffusion model for SE (RDM-SE) that explicitly incorporates deterministic repair into the encoder of diffusion model to address distortions caused by deterministic models. Different from GP-Unified [19], RDM-SE repairs deterministic information from only a subset of the encoder layers rather than all encoder layers and the shared bottleneck module. Recent studies suggest that feature downsampling may hinder further improvement of deterministic SE performance [11, 28, 29]. Thus, we adopt different downsampling strategies across the encoder layers to better repair deterministic information and extract diffusion embeddings. In the shallow encoder layers, we perform downsampling only along the frequency axis to better repair the distorted deterministic information. After getting the repaired deterministic-diffusion embedding, the diffusion-specific layers perform downsampling along the time axis or the time-frequency axis to extract embeddings better suited for score estimation. The choice between deterministic-only and deterministic-noisy conditions varies across deterministic models. We further enhance the RDM-SE with dual-streaming encoding to more effectively utilize both deterministic-only and deterministic-noisy conditions.
@@
Besides, although fine-grained processing deterministic models, such as TF-GridNet [29], demonstrate strong enhancement performance on simulated data, diffusion models with fine-grained enhanced deterministic conditions show less stable performance on real speech compared to the coarse-grained enhancement deterministic models with a UNet structure, such as DCCRN [27]. Thus, we propose a granularity progressive deterministic model, called COarse Fisrt then Fine EnhancEment (COFFEE), that combines fine-grained processing with the UNet structure. It enhances the feature starting from the bottleneck embedding, progressing to the reconstructed feature, and finally to the time-frequency bin using fine-grained processing.
@@ II-A Deterministic Methods for Speech Enhancement
Recent studies suggest that feature downsampling may hinder further improvement of SE performance [11, 28, 29]. The TasNet-series [11, 28] have attracted much attention due to their excellent separation ability. They use a separator to process the pseudo-spectrogram extracted by the Conv1d layer. In neither the Conv-TasNet with the temporal convolutional network (TCN) or dual-path RNN (DPRNN), features are rarely downsampled in the separator blocks. Some related works also show that extracting features with smaller strides performs better. With time-frequency (T-F) bin-level enhancement, TF-GridNet [29] shows more powerful performance and achieves state-of-the-art performance in many commonly used datasets. It does not do any downsampling processing. It first extracts the complex spectrogram; then, each T-F bin is enhanced twice along the time-domain and frequency-domain axes with several TF-GridNet blocks. With the small granularity processing, its processing speed is extremely slow.
TF-GridNet performs complex spectral mapping. The stacked real and imaginary (RI) parts of the complex spectrogram are used as input features. Each T-F bin first extracts a $D$-dimensional embedding with a Conv2D (kernel size $=3\times 3$ ) layer. The extracted feature is $R_{a}\in\mathbb{R}^{D\times T\times F}$, $T$ and $F$ represent the spectrogram’s time and frequency, respectively. Then, several TF-GridNet blocks enhance the noisy embeddings. An intra-frame full-band module, a sub-band temporal module, and a cross-frame self-attention module form the TF-GridNet block. The complex spectrogram embedding is first zero-padded to the shape $R_{b}\in D\times T\times F^{\prime}$ to ensure that adjacent information can be obtained when processing each T-F bin, where $F^{\prime}=\left\lfloor\frac{F-I}{J}\right\rfloor\times J+I$. $I$ and $J$ represent the kernel size and stride, respectively. In the intra-frame full-band module, $R_{b}$ is first unfolded as $R_{p}$:
@@
[…]
@@
In the cross-frame self-attention module, the frame embeddings are first computed. Then, the full-utterance self-attention is used to model long-range context information. The self-attention module has $L$ heads. The point-wise Conv2D, PReLU, layer normalization along the channel and frequency dimensions (cfLN) are used in each head to obtain query, key, and value. Then, the query, key, and value are reshaped as $Q_{l}\in\mathbb{R}^{T\times(F\times E)}$, $K_{l}\in\mathbb{R}^{T\times(F\times E)}$ and $V_{l}\in\mathbb{R}^{T\times(F\times D/L)}$. The point-wise Conv2D channels for computing the query and key are $E$, while the value is $D/L$. All three point-wise Conv2D layers have $D$ input channels. The attention outputs of all heads along the second dimension are concatenated and reshaped back to $D\times T\times F$. Another combination of the point-wise Conv2D with $D$ input and $D$ output channels followed by a PReLU and a cfLN is applied to aggregate cross-head information. Finally, the residual connection of the $Z_{b}$ is added to the output of the attention module.
After several TF-GridNet blocks processing, a 2D deconvolution (DeConv2D) with a $3\times 3$ kernel followed by linear units is used to obtain the predicted RI components of the enhanced speech and inverse short-time Fourier transform (iSTFT) is applied for signal re-synthesis.
@@ II-B Score-based Diffusion Model for Speech Enhancement
Diffusion model [31, 32, 33, 34] belongs to the probabilistic method [35]. The diffusion model has two processes: the forward process and the reverse process. The forward process follows the stochastic differential equation (SDE). A clean speech feature $\mathbf{x}_{0}$ is gradually transformed to a noisy speech $\mathbf{y}$ and white Gaussian noise [17, 22] with linear SDE:
@@
where $\mathbf{x}_{t}$ is the current state, and $t$ is a continuous state index $[0,T]$. $\mathrm{\mathbf{w}}$ represents a standard Wiener process, which is also called Brownian motion. $\mathbf{f}(\mathbf{x}_{t},\mathbf{y})$ is the drift coefficient, while $g(t)$ is the diffusion coefficient. The forward SDE has a corresponding reverse SDE, which is used to achieve SE [36, 33]. The score $\nabla_{\mathbf{x}_{t}}\mathrm{log}p_{t}(\mathbf{x}_{t}|\mathbf{y})$ on a state of the SDE is needed to be estimated for reverse SDE [36, 34]. A neural network $s_{\theta}$ with parameter set $\theta$ called the scoring model is used to approximate the score. The noisy speech $\mathbf{y}$ is used as the condition and makes the conditional probability as $p_{t}(\mathbf{x}_{t}|\mathbf{y})$. Thus, the inputs of the score models are defined as $s_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$. The reverse SDE is defined as follows:
@@
$\mathbf{x}_{t}\sim\mathcal{N}_{c}(\mathbf{x}_{t};\mathbf{y},\sigma(t)^{2}\mathbf{I}).$ (7)
@@
where $\mathcal{N}_{c}$ represents the circularly-symmetric complex normal distribution. $\mathbf{I}$ represents the identity matrix. $\sigma(t)^{2}$ is the variance of the white Gaussian noise included in $\mathbf{x}_{t}$ according to the forward SDE.
The training loss is based on the Mean Square Error (MSE) between true and estimated scores. The overall training objectives for denoising and TSE are derived as follows:

### v2

@@ Abstract
Score-based diffusion models for speech enhancement (SE) need to incorporate correct prior knowledge as reliable conditions to generate accurate predictions. However, providing reliable conditions using noisy features is challenging. One solution is to use features enhanced by deterministic methods as conditions. However, the information distortion and loss caused by deterministic methods might affect the diffusion process. In this paper, we first investigate the effects of using different deterministic SE models as conditions for diffusion. We validate two conditions depending on whether the noisy feature was used as part of the condition: one using only the deterministic feature (deterministic-only), and the other using both deterministic and noisy features (deterministic-noisy). Preliminary investigation found that using deterministic enhanced conditions improves DNSMOS and UTMOS on real data, while the choice between using deterministic-only or deterministic-noisy conditions depends on the deterministic models. Based on these findings, we propose the deterministic-diffusion unified model for SE to more effectively utilize both conditions. Moreover, we found that fine-grained deterministic models have greater potential in objective evaluation metrics, while UNet-based deterministic models provide more stable diffusion performance. Therefore, we also propose a deterministic model that combines coarse- and fine-grained processing for the diffusion. Experimental results on CHiME4 show that the proposed models effectively leverage deterministic models to achieve better SE evaluation scores on the DNSMOS and UTMOS for real evaluation sets. In addition, the proposed deterministic model proves to be more stable than other deterministic models when it is used for diffusion.
@@ I Introduction
Speech enhancement (SE) is commonly used to recover speech components from noisy speech [1, 2, 3, 4, 5] to improve the speech quality and intelligibility. Supervised SE methods [6] generally outperform traditional SE methods [1, 7] with sufficient training data available. These systems can be broadly categorized as either deterministic [8, 9, 10, 11, 12] or probabilistic [13, 14, 15, 16, 17, 18] approaches. Deterministic SE models learn an optimal deterministic mapping from noisy speech to clean speech [15], while probabilistic SE models capture the target distribution in either an implicit or explicit way [13, 14, 15, 16]. The presence of diverse and uncontrollable sources often makes the noise unpredictable, resulting in random signals. However, deterministic SE models only learn a deterministic mapping function, leading to performance degradation in unseen noise scenarios.
@@
Diffusion model [19, 20, 15, 14, 18, 21] is a probabilistic method learning a stochastic differential equation (SDE). It contains a forward diffusion process and a reverse diffusion process by following the SDE. The forward diffusion process gradually transforms the clean speech feature into noise, during which the neural network learns to reverse the incremental process of noise addition. The reverse process recovers the speech feature from noise with numerous reverse iterations. Different from the deterministic models [8, 9, 10, 11], SDE has the characteristics of the partially stochastic as: random noise is injected into each reverse SDE step. Score-based diffusion model for SE [14, 15, 22] takes the clean speech feature as the initial state and transforms it to the noisy speech feature with Gaussian white noise. Noisy speech feature is used as condition [14, 15, 22] and drift term [15, 22] to enable the model to capture noise information implicitly. However, noisy features pose a challenge in providing accurate prior knowledge as reliable conditions for diffusion. Schrödinger Bridge-based diffusion model learns the diffusion process directly between the noisy input and the clean distribution, rather than relying on Gaussian reference distributions, and it also achieves high performance in SE [23].
@@
Combining deterministic and diffusion models [24, 22, 17, 25] makes a promising solution for alleviating this challenge. Both noisy and deterministically enhanced (deterministic-noisy) features [25] or only deterministically enhanced (deterministic-only) features are used as conditions for diffusion [22]. When using a diffusion model to process deterministic enhanced features, the posterior samples are more likely to lie in high-density regions of the posterior probability for clean speech features given noisy speech features[22]. The effectiveness of the combination has been demonstrated [22, 24, 25], but no previous work has investigated on how different deterministic models affect the SE performance based on this pipeline.
@@
In this paper, we investigate the effects of using different deterministic SE models [26, 9, 27, 28] as conditions for diffusion-based SE. We experiment with score-based diffusion models that use either deterministic-only enhanced features (Deter-Diffusion) or deterministic-noisy features (StoRM [22] and GP-Unified [17]) as conditions. Deter-Diffusion and StoRM train the diffusion model only based on the score-matching loss, without incorporating any deterministic loss. Some studies have shown that deterministic information and diffusion-based information are complementary [20, 17]. Incorporating deterministic information—either from the encoder [17] or the decoder [20]—into diffusion models enhances performance. Rather than relying on complex fusion operations in the decoder [20], simply introducing the deterministic loss after the encoder during training also yields significant performance improvements [17]. GP-Unified [17] adopts the same encoder-decoder architecture as StoRM [22]. In addition to the score error, GP-Unified [17] introduces deterministic information into the diffusion model using an additional decoder. Although GP-Unified [17] outperforms StoRM [22], performing both deterministic estimation and diffusion within a shared bottleneck processing module may lead to performance degradation under noise-mismatch conditions [21], as both the deterministic and score-estimation information exhibit mismatched data distributions between simulated and real data.
@@
To more effectively utilize deterministic information as conditioning during diffusion, we propose a Deterministic-Diffusion Unified model for SE (D2U-SE), which explicitly incorporates the deterministic loss during the training of the diffusion model. The key difference from GP-Unified [17] is that D2U-SE incorporates deterministic information from only a subset of the encoder layers rather than all encoder layers and the shared bottleneck module. Recent studies suggest that feature downsampling may hinder further improvement of deterministic SE performance [9, 27, 28]. Thus, we adopt different downsampling strategies across the encoder layers to better incorporate deterministic information and extract diffusion embeddings. In the shallow encoder layers, we perform downsampling only along the frequency axis to better extract deterministic information. After getting the deterministic-diffusion embedding, the diffusion-specific layers perform downsampling along the time axis or the time-frequency axis to extract embeddings better suited for score estimation. Fig. 1 shows the high-level flowchart of the D2U-SE. The choice between deterministic-only and deterministic-noisy conditions varies across deterministic models. We further enhance the D2U-SE with Dual-streaming encoding (D3U-SE) to more effectively utilize both deterministic-only and deterministic-noisy conditions.
@@
Although fine-grained processing deterministic models, such as TF-GridNet [28], demonstrate strong enhancement performance on simulated data, diffusion models with fine-grained enhanced deterministic conditions show less stable performance on real speech compared to the coarse-grained enhancement deterministic models with a UNet structure, such as DCCRN [26]. Thus, we propose a granularity progressive deterministic model, called COarse Fisrt then Fine EnhancEment (COFFEE), that combines fine-grained processing with the UNet structure. It enhances the feature starting from the bottleneck embedding, progressing to the reconstructed feature, and finally to the time-frequency bin using fine-grained processing.
Figure 1: High-level conceptual flowchart of the D2U-SE framework. Each stage corresponds to a core modeling motivation.
@@ II-A Deterministic Methods for Speech Enhancement
Recent studies suggest that feature downsampling may hinder further improvement of SE performance [9, 27, 28]. The TasNet-series [9, 27] have attracted much attention due to their excellent separation ability. They use a separator to process the pseudo-spectrogram extracted by the Conv1d layer. In neither the Conv-TasNet with the temporal convolutional network (TCN) nor the dual-path RNN (DPRNN) are the features rarely downsampled in the separator blocks. Some related works also show that extracting features with smaller strides performs better. With time-frequency (T-F) bin-level enhancement, TF-GridNet [28] shows more powerful performance and achieves state-of-the-art performance in many commonly used datasets. It does not do any downsampling processing. It first extracts the complex spectrogram; then, each T-F bin is enhanced twice along the time-domain and frequency-domain axes with several TF-GridNet blocks. With the small granularity processing, the processing speed is extremely slow.
TF-GridNet performs complex spectral mapping. The stacked real and imaginary (RI) parts of the complex noisy spectrogram $\mathbf{Y}=\mathbf{Y}_{r}+j\mathbf{Y}_{i}\in\mathbb{C}^{T\times F}$ are used as input features, where $T$ and $F$ represent the spectrogram’s time and frequency. $j$ is the imaginary unit as $j^{2}=-1$. Each T-F bin first extracts a $D$-dimensional embedding with a Conv2D (kernel size $3\times 3$ ) layer. The extracted feature is $R_{a}\in\mathbb{R}^{D\times T\times F}$. Then, several TF-GridNet blocks enhance the noisy embeddings. An intra-frame full-band module, a sub-band temporal module, and a cross-frame self-attention module form the TF-GridNet block. The complex spectrogram embedding is first zero-padded to the shape $R_{b}\in D\times T\times F^{\prime}$ to ensure that adjacent information can be obtained when processing each T-F bin, where $F^{\prime}=\left\lfloor\frac{F-I}{J}\right\rfloor\times J+I$. $I$ and $J$ represent the kernel size and stride, respectively. In the intra-frame full-band module, $R_{b}$ is first unfolded as $R_{p}$:
@@
$R_{d}$ $=\left[\text{DeConv1D}(R_{t}[:,t,:]),\right.\left.\text{for }t=1,\ldots,T\right]$ (3)
@@
In the cross-frame self-attention module, the frame embeddings are first computed. Then, the full-utterance self-attention is used to model long-range context information. The self-attention module has $L$ heads. The point-wise Conv2D, PReLU, and layer normalization along the channel and frequency dimensions (cfLN) are used in each head to obtain query, key, and value. Then, the query, key, and value are reshaped as $Q_{l}\in\mathbb{R}^{T\times(F\times E)}$, $K_{l}\in\mathbb{R}^{T\times(F\times E)}$ and $V_{l}\in\mathbb{R}^{T\times(F\times D/L)}$. The point-wise Conv2D channels for computing the query and key are $E$, while the value is $D/L$. All three point-wise Conv2D layers have $D$ input channels. The attention outputs of all heads along the second dimension are concatenated and reshaped back to $D\times T\times F$. Another combination of the point-wise Conv2D with $D$ input and $D$ output channels, followed by a PReLU and a cfLN, is applied to aggregate cross-head information. Finally, the residual connection of the $Z_{b}$ is added to the output of the attention module.
After several TF-GridNet blocks processing, a 2D deconvolution (DeConv2D) with a $3\times 3$ kernel followed by linear units is used to obtain the predicted RI components of the enhanced speech $\tilde{\mathbf{X}}_{0}=\tilde{\mathbf{X}}_{0r}+j\tilde{\mathbf{X}}_{0i}\in\mathbb{C}^{T\times F}$. The training loss is based on the Mean Squared Error (MSE) between the clean and enhanced spectrogram of their stacked RI parts:
$\mathcal{L}_{Deter}=\left\|\tilde{\mathbf{X}}_{0}-\mathbf{X}_{0}\right\|^{2}_{2}$ (5)
where ${\mathbf{X}}_{0}={\mathbf{X}}_{0r}+j{\mathbf{X}}_{0i}\in\mathbb{C}^{T\times F}$ represents the clean spectrogram.
@@ II-B Score-based Diffusion Model for Speech Enhancement
Diffusion model [30, 31, 32, 33] belongs to the probabilistic method [34]. The diffusion model has two processes: the forward process and the reverse process. The forward process follows the stochastic differential equation (SDE). A clean speech feature $\mathbf{x}_{0}$ is gradually transformed to a noisy speech feature $\mathbf{y}$ and white Gaussian noise [15, 19] with linear SDE:
@@
where $\mathbf{X}_{t}$ is the current state, and $t$ is a continuous state index $[0,T]$. $\mathrm{\mathbf{w}}$ represents a standard Wiener process, which is also called Brownian motion. $\mathbf{f}(\mathbf{X}_{t},\mathbf{Y})$ is the drift coefficient, while $g(t)$ is the diffusion coefficient. The forward SDE has a corresponding reverse SDE, which is used to achieve SE [35, 32]. The score $\nabla_{\mathbf{X}_{t}}\mathrm{log}p_{t}(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{Y})$ on a state of the SDE needs to be estimated for reverse SDE [35, 33]. A neural network $s_{\theta}$ with parameter set $\theta$, called the scoring model, is used to approximate the score. The noisy speech feature $\mathbf{Y}$ is used as the condition and makes the conditional probability as $p_{t}(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{Y})$. Thus, the inputs of the score models are defined as $s_{\theta}(\mathbf{X}_{t},\mathbf{Y},t)$. The reverse SDE is defined as follows:
@@
$\mathbf{X}_{t}\sim\mathcal{N}_{c}(\mathbf{X}_{t};\mathbf{Y},\sigma(t)^{2}\mathbf{I}),$ (8)
@@
where $\mathcal{N}_{c}$ represents the circularly-symmetric complex normal distribution. $\mathbf{I}$ represents the identity matrix. $\sigma(t)^{2}$ is the variance of the white Gaussian noise included in $\mathbf{X}_{t}$ according to the forward SDE:
$\sigma(t)^{2}=\frac{c(k^{2t}-e^{-2\gamma t})}{2(\gamma+\log k)},$ (9)
where $k$, $\gamma$, and $c$ are positive scalar constants.
The overall training objectives for denoising are derived as follows:
$\underset{\theta}{\mathrm{arg}~\mathrm{min}\,}\mathbb{E}\left[\left\|\mathrm{s}_{\theta}+\mathrm{z}/{\sigma(t)}\right\|^{2}_{2}\right],$ (10)
where $\mathrm{z}\sim\mathcal{N_{C}}(\mathrm{z};0,\mathbf{I})$ is a sampled white Gaussian noise with an identity matrix $\mathbf{I}$. The loss function is defined as denoising score matching:
$\mathcal{L}_{\mathrm{Score}}=\lambda(t)\left\|s_{\theta}(\mathbf{X}_{t},\mathbf{Y},t)-\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{X}_{0},\mathbf{Y})\right\|_{2}^{2},$ (11)
where $\lambda(t)$ is the weighting function that depends on $t$.
II-C The Relationship Between Score-based Diffusion Model and Deterministic Model for Speech Enhancement
Based on the formulation in Eq. (11), denoising score matching can be expressed as training a denoiser model $D_{\theta}$ to match the clean signal estimate by [36, 37]:
$\mathcal{L}_{\text{Score-denoiser}}=\lambda(t)\left\|D_{\theta}(\mathbf{X}_{t},\mathbf{Y},t)-\mu_{t}(\mathbf{X}_{0},\mathbf{Y})\right\|_{2}^{2}$ (12)
$=\lambda(t)\left\|D_{\theta}(\mathbf{X}_{t},\mathbf{Y},t)-(1-e^{-\gamma t})\mathbf{Y}-e^{-\gamma t}\mathbf{X}_{0}\right\|_{2}^{2}.$
When $t=0$,
$\mathcal{L}_{\text{Score-denoiser}}=\lambda(t)\left\|D_{\theta}(\mathbf{X}_{t},\mathbf{Y},t)-\mathbf{X}_{0}\right\|_{2}^{2}.$ (13)
Similarly, the training loss of the deterministic model can also be expressed in a similar format based on Eq. (5):
$\mathcal{L}_{\text{Deter-denoiser}}=\left\|D_{\theta}(\mathbf{Y})-\mathbf{X}_{0}\right\|_{2}^{2}.$ (14)
Compared with Eq. (13) and Eq. (14), both the deterministic and score-based diffusion models aim to recover the clean feature from its noisy feature. The key differences lie in the presence of the time-dependent scaling factor and in whether the noisy feature $\mathbf{Y}$ is explicitly incorporated during the denoising process. Besides, the choice of the weighting function $\lambda(t)$ can have a substantial impact in practice and may lead to slight performance differences.
III Proposed Method
We propose combining the deterministic and diffusion models for SE. It contains a deterministic model and a diffusion model as shown in Fig. 2. In the deterministic model, we propose a granularity progressive deterministic model COFFEE by leveraging the advantages of both high-speed down-sampling processing and high-performance fine-granularity processing as shown in Fig. 2. The bottleneck embedding, reconstructed features in the decoder, and the T-F bins are enhanced with fine-granularity processing gradually.
To better integrate the deterministic and diffusion models, we propose the D2U-SE. We design the encoder of D2U-SE based on the characteristics of deterministic and diffusion processes. Within the encoder, the features are processed through several deterministic-diffusion layers and diffusion-specific layers. Deterministic information is explicitly introduced into the shared encoder layers by computing the deterministic loss. The deterministic-diffusion shared layers perform downsampling only along the frequency axis to retain sufficient information for learning deterministic information. The diffusion-specific layers focus on extracting embeddings only for diffusion. Besides, the deterministic models affect the choice between using deterministic-only and deterministic-noisy conditions. Thus, we further improve the D2U-SE with dual-stream encoding (D3U-SE) to simultaneously utilize both deterministic-only and deterministic-noisy conditions.
Figure 2: Flowchart of the proposed method. It includes the proposed COFFEE deterministic model and the proposed D2U-SE. Within COFFEE, we introduce Time-Embedding (T-E) bin-level, Time-Reconstructed Feature (T-RF) bin-level, and Time-Frequency (T-F) bin-level processing. Within D2U-SE, the skip connections between the encoder and decoder are omitted in the figure for clarity. Lowercase letters denote the time-domain representations of their corresponding uppercase symbols.
III-A Deterministic Method: COFFEE
We use a complex-valued network, which often requires decomposing a signal into its real and imaginary components, potentially avoiding the loss of important phase information and enhancing performance in certain tasks [38]. It contains an encoder, a decoder, and the bottleneck processing layers. The encoder reduces the resolution of the input feature and extracts high-level representations. The encoder layer processes the input complex feature $\mathbf{Y}=\mathbf{Y}_{r}+j\mathbf{Y}_{i}\in\mathbb{C}^{T\times F}$ with complex Conv2d. The complex-valued filter $W=W_{r}+jW_{i}$ is used for convolution operation:
$\mathbf{Y}_{o}=(\mathbf{Y}_{r}*W_{r}^{e}-\mathbf{Y}_{i}*W_{i}^{e})+j(\mathbf{Y}_{r}*W_{i}^{e}+\mathbf{Y}_{i}*W_{r}^{e}),$ (15)
The real-valued matrices $W_{r}^{e}$ and $W_{i}^{e}$ represent the real and imaginary parts of the complex convolution kernel, respectively. $W_{r}^{e}$ and $W_{i}^{e}$ are implemented with the Conv2d layer with different parameters. $\mathbf{Y}_{o}$ is the filtered feature. Batch normalization and PReLU follow the Conv2d layers.